prove that cosx>=1-(x^2/2)for all x in the interval[0,2π]
1
Expert's answer
2020-09-07T17:16:58-0400

"Consider \;the \;following\;function\;:
\\
f(x)= \cos(x) -1 +\dfrac{x^2}{2}
\\
Then\;the\;assertion\;that\; \cos(x) \geq 1 -\dfrac{x^2}{2}\; is \;identical \;to\; showing \;that\;:
\\
\;\;\;\;\;\;\;\;\;\;\cos(x) -1 +\dfrac{x^2}{2} \geq0
\iff f(x)\geq0"


"When \;
x
=
0,\;
 we\; have\; 
f
(
0
)
=
0."


"So\;to\;prove\;that\;f(x) \geq 0 \;we\; attempt \;to \;verify\; that\; the\; function\;f(x)
\\is \;strictly\;increasing.To \;do \;this, \;take \;the \;derivative:"

"f'(x)=-\sin(x)+x"


"If \;we\; show 
f
'
(
x
)
 \;is\; strictly\; positive\; for\; x>0,

 then \;we\; can\; apply\; the \;mean \;value\;
\\
 theorem \;to\; show\; 

f
(
x
)\;
 must\; be \;strictly\; increasing. \;Differentiating \;again\; we\; get:"


"f''(x)=-\cos(x)+1"


"Clearly , \;as \;\cos(x)\leq 1 \implies 1-\cos(x) \geq0 \;and\;so\;we\;can\;use\;this
\\
(and\;mean\;value\;theorem\;again) \;to\;prove\;f'(x)>0\; \forall \;x\isin[0,2\pi],\;hence\;the\;result
\\
follows."